# Perturbative Chern-Simons invariants from non-acyclic flat connections

### v1

@@ 5. Discussion
One could ask if there are other (potentially simpler) ways to generalize the perturbative Chern-Simons invariants to non-acyclic flat connections. One thing one can do is: fix a non-acyclic flat connection $A$ and a metric $g$, and consider the “BV effective action on zero modes” – the form degree zero part of $\check{W}$ introduced in equation (4), evaluated at $(A,A,g)$. One can show ([7],[12]) that it is independent of the metric, up to a $\Delta_{\mathsf{a}}$-exact term, where $\Delta_{\mathsf{a}}$ denotes the BV Laplacian on formal half-densities. However, if $\dim H^{1}_{A}>0$, then the relevant BV cohomology is trivial. [^20] The effective action has ghost number 0, and so corresponds (under the odd Fourier transform) to a top form. The cohomology of formal differential forms however, by Poincaré Lemma, is concentrated in degree 0 (corresponding to ghost degree $\dim H^{1}_{A}$).
Another way would be to consider the top form on $\mathcal{M}^{\prime}$ given by the taking top form component of $\overline{Y}$ and setting the fiber coordinates to 0. This amounts to taking only diagrams appearing in the Axelrod-Singer construction, considering them in a family over the reference flat connection $A$, and multiplying them with the square root of the Ray-Singer torsion. In the lowest degree, this means considering only diagrams $\Gamma_{1}$ and $\Gamma_{2}$ in Figure 1. At this order, one can check (29) explicitly by a diagrammatic computation. In this computation, one term in the metric variation of $\Phi_{\Gamma_{1}}+\Phi_{\Gamma_{2}}$ is not cancelled by the de Rham differential of the 1-loop terms. The remaining term, however, is cancelled in the metric variation of $\Phi_{\Gamma_{3}}$. This suggests that the variation of the top-form including only the Axelrod-Singer diagrams with respect to the metric is not exact, and thus its cohomology class does not define an invariant. Only after summing over all diagrams in Figure 1 does one obtain a metric invariance up to exact terms. One reason these corrections arise is the fact that to obtain a global, equivariant family of chain contractions $(i_{A},p_{A},K_{A})$, we need to vary the “gauge-fixing operator” $d_{A}^{*}$ together with the flat connection $A$ we are expanding around.
@@ Appendix A Formal Geometry on the moduli space of flat connections
In an attempt at self-containedness, we review certain aspects of formal geometry on the moduli space of flat connections. Formal geometry has been used extensively in the context of perturbative BV quantization, see [6], [4]. On the moduli space of flat connections it was considered in [8, Section 2].

### v2

@@ 5. Discussion
One could ask if there are other (potentially simpler) ways to generalize the perturbative Chern-Simons invariants to non-acyclic flat connections. One thing one can do is: fix a non-acyclic flat connection $A$ and a metric $g$, and consider the “BV effective action on zero modes” – the form degree zero part of $\check{W}$ introduced in equation (16), evaluated at $(A,A,g)$. One can show ([7],[13]) that it is independent of the metric, up to a $\Delta_{\mathsf{a}}$-exact term, where $\Delta_{\mathsf{a}}$ denotes the BV Laplacian on formal half-densities, and thus defines an invariant. However, if $\dim H^{1}_{A}>0$, then the relevant BV cohomology, and therefore also this invariant, is trivial. [^26] The effective action has ghost number 0, and so corresponds (under the odd Fourier transform) to a top form. The cohomology of formal differential forms however, by Poincaré Lemma, is concentrated in degree 0 (corresponding to ghost degree $\dim H^{1}_{A}$).  [^27] A side remark: In [7, Proposition 6] one has a situation – the effective BV action of Chern-Simons theory on a rational homology sphere computed at zero connection – where the BV cohomology is trivial, but one can define a certain refinement of BV cohomology, quotienting BV cocycles by a restricted class of coboundaries coming from changes of gauge-fixing in the path integral. This refinement is nontrivial, and the refined BV cohomology class of the effective action is a meaningful invariant. However, it is not clear how to define such a refinement for irreducible smooth connections.
Another way would be to consider the top form on $\mathcal{M}^{\prime}$ given by taking the zero form component of $\overline{Y}$, setting the fiber coordinates to zero and interpreting the result as a density/top form on $\mathcal{M}^{\prime}$. This amounts to taking only diagrams appearing in the Axelrod-Singer construction, considering them in a family over the reference flat connection $A$, and multiplying them with the square root of the Ray-Singer torsion. In the lowest degree, this means considering only diagrams $\Gamma_{1}$ and $\Gamma_{2}$ in Figure 1. At this order, one can check (30) explicitly by a diagrammatic computation, see Appendix B. In this computation, one term in the metric variation of $\Phi_{\Gamma_{1}}+\Phi_{\Gamma_{2}}$ is not cancelled by the de Rham differential of the 1-loop terms. The remaining term, however, is cancelled in the metric variation of $\Phi_{\Gamma_{3}}$, cf. cancellation $(1,3)+(3,4)+(7,6)=0$ in (44). This suggests that the variation of the top-form including only the Axelrod-Singer diagrams with respect to the metric is not exact, and thus its cohomology class does not define an invariant. Only after summing over all diagrams in Figure 1 does one obtain a metric invariance up to exact terms. One reason these corrections arise is the fact that to obtain a global, equivariant family of chain contractions $(i_{A},p_{A},K_{A})$, we need to vary the “gauge-fixing operator” $d_{A}^{*}$ together with the flat connection $A$ we are expanding around.
@@ Appendix A Formal Geometry on the moduli space of flat connections
In an attempt at self-containedness, we review certain aspects of formal geometry on the moduli space of flat connections. Formal geometry has been used extensively in the context of perturbative BV quantization, see [6], [4]. On the moduli space of flat connections it was considered in [9, Section 2].
@@ A.2.2. A formula for the Grothendieck connection of $\underline{\varphi}^{g}$
Appendix B A sketch of direct diagrammatic proof of metric independence of $\Lambda_{1}$
Metric independence of the invariant in degree $O(\epsilon^{1})$, $\Lambda_{1}$ (14), is of course a part of Theorem 2.1, but one can also check it directly. We sketch the argument here (modulo signs), as a sanity check.
Denote
(42) $\lambda=D_{1}+D_{2}+D_{3}+D_{4}+D_{5}+D_{6}+D_{\mathrm{grav}}$
the integrand in (14), with $D_{i}=\tau_{A}^{\frac{1}{2}}\Phi_{i}$ for $i=1,\ldots,6$ and $D_{\mathrm{grav}}=-\frac{\langle f,f\rangle}{24(2\pi)^{2}}\tau_{A}^{\frac{1}{2}}S_{\mathrm{grav}}(g,\phi)$. Also let $\Xi=D_{7}+D_{8}+D_{9}+D_{10}$ be as in (15). (Recall that $\lambda$, $\Xi$ are the $O(\epsilon^{1})$ terms in $\check{Y}^{\mathrm{glob},(0)}$ and $\check{Y}^{\mathrm{glob},(1)}$). We want to prove that
(43) $\delta_{g}\lambda+d_{\mathcal{M}^{\prime}}\Xi=0.$
One can do it directly by applying $\delta_{g}$ to terms in $\lambda$ and applying $d_{\mathcal{M}^{\prime}}$ to terms in $\Xi$, write the results as sums of (weights of) certain Feynman diagrams – see Figure 3 – and then track the cancellations of diagrams.
Figure 3. Diagrammatic formulae for $\delta_{g}$ applied to terms in $\lambda$ and $d_{\mathcal{M}^{\prime}}$ applied to terms in $\Xi$.
In Figure 3, for $i=1,\ldots,6$, we label the diagrams computing $\delta_{g}D_{i}$ by $(i,j)$. Likewise, for $i=7,\ldots,10$, we label the diagrams computing $d_{\mathcal{M}^{\prime}}D_{i}$ by $(i,j)$.
This computation is based on identities
$\delta_{g}K$ $=$ $-[d,KH_{\delta g}K]-KH_{\delta g}P-PH_{\delta g}K,$
$\delta_{g}i$ $=$ $-dKH_{\delta g}i,$
$\mathcal{L}_{\chi}K$ $=$ $-[d,KH_{\delta A^{\prime}=\chi}K]+KH_{\delta A^{\prime}=\chi}P-PH_{\delta A^{\prime}=\chi}K-K\mathrm{ad}_{\chi}K,$
$\mathcal{L}_{\chi}i$ $=$ $-dKH_{\delta A^{\prime}=\chi}i-K\mathrm{ad}_{\chi}i,$
$\mathcal{L}_{\chi}\tau$ $=$ $\tau\cdot\mathrm{Str}\,K\mathrm{ad}_{\chi}.$
Here $\chi\in\mathrm{Harm}^{1}_{A}\cong T_{A}\mathcal{M}^{\prime}$ and $\mathcal{L}_{\chi}$ is the directional derivative in the direction $\chi$ on the moduli space; when acting on the torsion $\tau$, $\mathcal{L}_{\chi}$ uses an appropriate connection in the determinant line bundle, see [9, Lemma D.1]. [^35] There are also some identities used for simplification, e.g., $KH_{\delta A^{\prime}=d\beta}K=K\mathrm{ad}_{\beta}K$, $KH_{\delta A^{\prime}=d\beta}P=K\mathrm{ad}_{\beta}P$, for $\beta\in\Omega^{0}(M,\mathfrak{g})$.
In Figure 3 we give labels $a,b,\ldots$ to elements of the Feynman graph on the left (we pick one label for each orbit of $\mathrm{Aut}(\Gamma)$) and on the right we indicate which terms arise from the derivative hitting which element of the graph on the left. [^36] Including the possibility that the derivative hits the torsion factor $\tau_{A}^{\frac{1}{2}}$ implicitly contained in $D_{i}$. This happens in terms $(7,9)$, $(8,7)$, $(9,7)$, $(10,7)$.
The conventions in Figure 3 (the assignment of tensors to edges and vertices – the Feynman rules) are as follows.
Edges: A solid edge is assigned $K$; a dashed edge is assigned the projection onto harmonic forms $P$; a solid-dashed edge is the projection onto harmonic 2-forms if the solid end is treated as “out” and onto harmonic 1-forms if the solid end is “in;” a solid double edge is the identity; [^37] Identity on the double edge $e=(uv)$ is the interpretation if the diagram is seen as an integral over $M^{\times V}$. Alternatively, in terms of the compactified configuration space $C_{V}(M)$, one considers the integral over the principal boundary stratum where $u$ and $v$ approach one another (and other propagators connecting $u$ and $v$ are replaced with their continuous part at $u=v$.) a solid edge ending with a square box is $dK$.
Vertices: An ordinary trivalent vertex is the wedge product of forms tensored with the Lie bracket in $\mathfrak{g}$; a white circle vertex is assigned $H_{\delta A^{\prime}=\chi}$ with $\chi\in\mathrm{Harm}^{1}_{A}$ the argument coming from the side half-edge; a black vertex is $H_{\delta g}$. A circled black vertex is $\mathcal{L}_{\chi}H_{\delta g}$ ($\chi$ decorates the side half-edge). A circled white vertex is $\delta_{g}H_{\delta A^{\prime}=\chi}$.
Additionally, each diagram is understood to contain the torsion factor $\tau_{A}^{\frac{1}{2}}$.
Cancellations: The term $\mathbb{A}$ in $\delta_{g}D_{1}$ is the anomalous part of the metric dependence of the Theta graph, as in [2]. It is cancelled by $\delta_{g}D_{\mathrm{grav}}$. Diagrams $(1,1)$ and $(2,1)$ cancel by Jacobi identity. The remaining diagrams cancel as follows:
(44) (1,2)+(8,3)=0;(1,3)+(3,4)+(7,6)=0;(2,2)+(8,7)=0;(2,3)+(8,5)=0;(2,4)+(4,5)+(7,3)=0;(2,5)+(7,7)=0;(2,6)+(7,9)=0;(3,1)+(8,1)=0;(3,2)+(10,4)=0;(3,3)+(7,4)=0;(3,5)+(8,6)+(5,1)+(9,3)=0;(4,1)+(10,2)=0;(4,2)+(10,7)=0;(4,3)+(8,4)=0;(4,4)+(6,1)+(10,1)=0;(4,6)+(7,8)=0;(5,2)+(9,6)=0;(6,2)+(10,6)=0;(7,1)+(9,2)=0;(7,2)+(9,7)=0;(7,5)+(9,1)+(10,3)=0;(8,2)+(9,5)=0;(9,4)+(10,5)=0.(1,2)+(8,3)=0;\;(1,3)+(3,4)+(7,6)=0;(2,2)+(8,7)=0;\\ (2,3)+(8,5)=0;\;(2,4)+(4,5)+(7,3)=0;\;(2,5)+(7,7)=0;\;(2,6)+(7,9)=0;\\ (3,1)+(8,1)=0;\;(3,2)+(10,4)=0;\;(3,3)+(7,4)=0;\;(3,5)+(8,6)+(5,1)+(9,3)=0;\\ (4,1)+(10,2)=0;\;(4,2)+(10,7)=0;\;(4,3)+(8,4)=0;\;(4,4)+(6,1)+(10,1)=0;\\ (4,6)+(7,8)=0;\;(5,2)+(9,6)=0;\;(6,2)+(10,6)=0;\;(7,1)+(9,2)=0;\\ (7,2)+(9,7)=0;\;(7,5)+(9,1)+(10,3)=0;\;(8,2)+(9,5)=0;\;(9,4)+(10,5)=0.
This proves (43).